OpenStudy (anonymous):

Solve the equation: 10x-9 = 4(x -1) + 1

OpenStudy (anonymous):

right side = (4x-4)+1 10x-9=4x-3 cancel -9 on the left side by adding to both sides 10x=4x+6 cancel 4x on the right side by subtracting 4x to both side 6x=6 divide and get x by itself x=1

OpenStudy (anonymous):

check by substituting x=1 back in 10((1)-9) = 4((1)-1)+1 1=1

OpenStudy (anonymous):

fosho thanks